y is between x and z. t is between y and z. if xz= 82, yt=22, and xy=tz. what is tz?
1
Expert's answer
2020-08-28T12:52:58-0400

xz = XY+YT+TZ


XY = TZ


Next,


82 = XY + 22 + TZ


XY +TZ = 82-22


XY + TZ =60


XY = TZ =60/2 = 30


XT = XY + YT = 30+ 22 =52


Answer: TZ =30.